If wrong will mark thumbs down, please please only answer if you know the answer is correct :)

Chapter 05, Problem 001 2 Your answer is partially correct. Try again. Only two horizontal forces act on a 7.5 kg body. One force is 5.4 N, acting due east, and the other is 9.6 N, acting 68° north of west. What is the magnitude of the bodys acceleration? Number 8.90320 units m/s^2C Number | 8.9032 m/s^ the tolerance is +/-2%
Answer #1

F1 = 5.4N E

F2 = 9.6N, 68 degrees N of W

Net horizontal force (i.e. along east):

Fx = F1 – F2 cos(68) = 1.8N

Net vertical force (i.e. along North):

Fy = F2sin(68) = 8.9 N

Net force:

F = sqrt (Fx^2 + Fy^2) = 9.08N

Thus,

acceleration a = F/m

= 9.08/7.5 = 1.21 m/s2